OpenStudy (anonymous):

Need help plz evaluate x divided by 80 for x=8/9

OpenStudy (anonymous):

A,. 71 1/9 B. 1/90 C. 9/640 D. 90

OpenStudy (anonymous):

i think its B or D

OpenStudy (anonymous):

Tell me how old you are so I can help you better.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

ok. tell me one thing. do you know what is 1 divided by 4?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

25??

OpenStudy (anonymous):

write it as a fraction.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

very good. So to evaluate x divided by 4 for x=1 we will get 1/4 you got it?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Be sure about it. :) what about to evaluate x divided by 2 for 1/3?

OpenStudy (anonymous):

U would do 1/2?

OpenStudy (anonymous):

2) Evaluate X divided by 3 for X=1/2 \[x \times \frac{ 1 }{ 3 }-->\frac{ 1 }{ 2 }\times\frac{ 1 }{ 3 }=\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

Do you understand what I did?

OpenStudy (anonymous):

You multiplied 1/2 and 1/3 then u got 1/6

OpenStudy (anonymous):

yes! I multiplied X by 1/3. That exactly means that I divided X by 3.

OpenStudy (anonymous):

So, what if I want to divide a number by 80? What should I do? I need to multiply it by which number?

OpenStudy (anonymous):

U need to multiply it by 1/8?

OpenStudy (anonymous):

Or by 1/2

OpenStudy (anonymous):

1/8? Didin't you forget something like a zero?

OpenStudy (anonymous):

You should multiply it by 1/80. I guess you just forgot to type the zero. What if you want to divide a number by 12?

OpenStudy (anonymous):

U would divide it by 1 1/2 maybe

OpenStudy (anonymous):

You know the answer. But you type it incorrectly. Type it correctly.

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

1/12!!!!

OpenStudy (anonymous):

what if you want to divide a number by 50?

OpenStudy (anonymous):

@Artteam12 please do it faster. I have to go to my class. I want you to finish it correctly. Answer me faster please.

OpenStudy (anonymous):

Its ok my sister said she would help merry christmas bye

OpenStudy (anonymous):

aaaaaaaaaaah, hahaha Ok! Have so much fun!

OpenStudy (anonymous):

I wish I could make you understand. Anyhow! For dividing a fraction by 80 you must multiply that fraction by 1/80! It's very easy. and here are 2 other examples of this.

OpenStudy (anonymous):

If u have to gget back to ur class get back to ur class i dont want to waist ur time

OpenStudy (anonymous):

Dividing X by 15 for X=5/3 --> 5/3*1/15=5/45=1/9 Dividing X by 40 for X=4/10 --> 4/10*1/40=4/400=1/100 You should simplify denominator and numerator. Do Great!